OpenStudy (anonymous):

A rectangular prism has length 15m , width 2m, and volume 180m3 . What is the height of the prism?

OpenStudy (lgbasallote):

\[V = \pi r^2 h\] therefore... \[h = \frac{V}{\pi r^2}\]

OpenStudy (lgbasallote):

does that answer your question??